A speeding car is travelling at a constant speed of 44 ms–1 when it passes a stationary police car. The police
car immediately accelerates uniformly from rest at a rate of 2.2 ms–2. If the car does not slow down and the
police officer maintains the rate of acceleration, how long will it take the police car to catch the speeding car?
1
Expert's answer
2019-09-11T10:44:18-0400

"v_1=44m\/s"

"a_1=0"

"a_2=2.2m\/s^2"

"S_1=S_2"

________________

"t-?"

________________

"S_1=vt"

"S_2=at^2\/2"

________________

"at^2\/2=vt"

"2.2t^2\/2=44t"

"2.2t^2\/2-44t=0"

"1.1t(t-40)=0"

"t_1=0"

"t_2=40"